Joe lives 5 miles from the school and Maya lives 2 miles from the school. If joe walks from his house in anopposite direction from the school at a speed of 4 miles per hour, and Maya walks from her house in an opposite direction from the school at a speed of 6 miles per hour, after how long will they be the same distance from the school? What is that distance?
Answer by shree840(260)

5+4t=2+6t
3=2t or t=3/2 ANS